OpenStudy (omnomnom):

A object is thrown straight upwards with a speed of 19.6 m/sec. Three seconds later, its speed is: Select one: a. zero. b. 9.8 m/sec downward. c. 49 m/sec upward. d. 29.4 m/sec downward.

OpenStudy (anonymous):

Okay, do you know the downward acceleration due to gravity?

OpenStudy (omnomnom):

Nope it doesnt say

OpenStudy (anonymous):

That is because you're supposed to memorize it. It is always \(9.80665 m/s^2\).

OpenStudy (anonymous):

You can round it to \(9.81m/s^2\) for this problem.

OpenStudy (anonymous):

So you want to change from \(\large \frac{m}{s^2}\) to \(\large \frac{m}{s}\)

OpenStudy (anonymous):

That's a hint that you usually want to multiply by \(s\).

OpenStudy (omnomnom):

oh ok

OpenStudy (anonymous):

If you accelerate at a constant rate for a period of time, the change in velocity is acceleration times time.

OpenStudy (anonymous):

So first, what is the change in velocity?

OpenStudy (omnomnom):

um... Vf-Vi

OpenStudy (anonymous):

Yes, that is true. But we want to calculate it by using acceleration and time, since that is what we know.

OpenStudy (omnomnom):

so vf=Vi+acceleration(time) \[Vf = Vi+ a (t)\]

OpenStudy (anonymous):

Yeah. We are given time and initial velocity (and we keep acceleration to memory).

OpenStudy (anonymous):

So we should be able to figure this out.

OpenStudy (omnomnom):

Yay :D

OpenStudy (omnomnom):

the answer i got wasnt in any of the choices

OpenStudy (anonymous):

Oh, well one thing to note is that gravity is downwards.

OpenStudy (omnomnom):

so its -9.8 m/s^2

OpenStudy (anonymous):

So it will be in the opposite direction of the initial velocity.

OpenStudy (anonymous):

That is one way of doing it.

OpenStudy (omnomnom):

? how would it be opposite?

OpenStudy (anonymous):

You can chose upward to be positive and downward to be negative. You can also choose upward to be negative and downward to be positive. As long as you are consistent.

OpenStudy (anonymous):

The initial velocity is upward, gravity is downward. They are in the opposite direction.

OpenStudy (omnomnom):

yeah but either way i got 19.6m/s+-88.2m

OpenStudy (omnomnom):

\[Vf = 19.6\frac{ m }{ s } +( -9.8 \frac{ m }{ s^2 } ) (3^2) \]

OpenStudy (anonymous):

That 3 shouldn't be squared!

OpenStudy (omnomnom):

Omg i just noticed THANK YOU